OpenStudy (anonymous):

Identify the resulting probability as theoretical or experimental. The outcomes for rolling a number cube with faces labeled 1− 6 are equally likely. Find the probability that the number rolled is at least a 4. A. The experimental probability is 1/2 . B. The experimental probability is 1/3 . C. The theoretical probability is 1/2 . D. The theoretical probability is 1/3 .

OpenStudy (anonymous):

I need help.........!!!!!!!!!!!

OpenStudy (k12awesomeness):

I'll help you.. Hold on.

OpenStudy (knov):

the cub with faces labeled one to six. Our desired outcome is four on a face of the cube.

OpenStudy (anonymous):

Uuuuhhhhhhh

OpenStudy (k12awesomeness):

It's C.

OpenStudy (knov):

let's assume we threw cub a 6 times.

OpenStudy (knov):

pls don't give out the answer.

OpenStudy (knov):

you can be reported.

OpenStudy (k12awesomeness):

whatever

OpenStudy (knov):

you think you helped him ?

OpenStudy (anonymous):

I'm actually here to learn. Not be given direct answers.

OpenStudy (k12awesomeness):

I was going to explain... you just didn't give me a chance.

OpenStudy (knov):

well, you could team up with me.

OpenStudy (k12awesomeness):

sure.

OpenStudy (knov):

anyways, back to the question.

OpenStudy (knov):

let's assume you throw cube a infinite times.

OpenStudy (knov):



OpenStudy (k12awesomeness):

theoretical means, the prob you got with out trials just basing on what you think to be true experimental is when you look at the data, and then say oh look i got 4, 2 times out of 8 times so 2/8 prob

OpenStudy (k12awesomeness):

so "at least" a 4, means it has to be a 4 or 5 or 6, that is 3 out of the 6 possibilities, theoretically thats 1/2

OpenStudy (k12awesomeness):

Does that make sense?

OpenStudy (anonymous):

Yes

OpenStudy (knov):

u sure?

OpenStudy (knov):

i mean @WWENXTFAN

OpenStudy (anonymous):

I'm 85% sure.

OpenStudy (knov):

let's assume we have a coin. and we throw it a 2 times.

OpenStudy (anonymous):

Ok

OpenStudy (knov):

what's the probability of it lands on heads.

OpenStudy (k12awesomeness):

There's also this way to explain it... By stating "at LEAST a 4", it means that the desired numbers are 4 or higher: 4, 5, 6 That's 3 numbers out of 6. This can be modeled by the fraction 3/6 This is also equal to 1/2. Hope that helps!

OpenStudy (knov):

remember that we only have 2 outcomes. tails or heads. you would see on heads or on tails.

OpenStudy (knov):

so it's 1/2.

OpenStudy (knov):

it's a coin. @K12awesomeness

OpenStudy (knov):

then what you meant make no sense to me.

OpenStudy (knov):

@WWENXTFAN i hope you understand it.

OpenStudy (anonymous):

Hold on I have to clear my mind. My youngest sister stresses me out a lot.

OpenStudy (k12awesomeness):

Your not the only on who's sister stresses you out....

OpenStudy (anonymous):

She is getting mad over a drink.

OpenStudy (k12awesomeness):

mines gettin mad over a toy!

OpenStudy (k12awesomeness):

anyway.. do you understand why it's C. ?

OpenStudy (anonymous):

I think so

OpenStudy (k12awesomeness):

okay, good.

OpenStudy (anonymous):

Ok thanks @K12awesomeness and @knov.

OpenStudy (k12awesomeness):

No Problem!

OpenStudy (anonymous):

Okay bye. :)

OpenStudy (k12awesomeness):

bye!

OpenStudy (k12awesomeness):

:)

OpenStudy (welshfella):

the experimental probability is the result of experiment. You might throw the cube say 100 times and record the numbers you get on each throw. from those numbers you can work out the experimental probability. The theoretical probability is what you would expect in theory. For example you would expect a probability of getting a 4 of 1/6 because there is one 4 and total of 6 numbers. Theoretical probability of getting at least 4 is 1/2 :- P(getting 4,5,6) = number of favourable results / number of possible results = 3/6 = 1/2